I have been going over this question for an hour now and i just can't seem to get it. Please help.
The x axis goes from –5 to 5 and the y axis goes from –3 to 7.
a) What is the slope of this line?
b) What is the horizontal intercept of this line?
c) What is the vertical intercept of this line?
d) Write an equation of the form y = mx + b for this linear function.

Answer by checkley71(8403)

IF YOU MEAN THAT THESE POINTS ARE (-5,-3)(5,7) THEN THE SLOPE IS
(7+3)/(5+5)=10/10=1
Y=mX+b
-3=1*-5+b
-3=-5+b
-3+5=b
b=2 OR THE Y INTERCEPT
SETTING Y=0 WE GET THE X INTERCEPT THUS
0=1X+2
X=-2 FOR THE X INTERCEPT
THE EQUATION FOR THIS LINE IS
Y=X+2